integrate x^1/2+x^-1/2 ysdfghjgfxgjhjjjjjjjjjjjjjjj sdfghjjdhhnb

Grade:11

1 Answers

sandeep
7 years ago
\int (x^{1/2}+x^{-1/2}) dx
\int (x^{1/2})dx+\int( x^{-1/2})dx
\frac{x^{(1/2)+1}}{(1/2)+1)} + \frac{x^{(-1/2)+1}}{(-1/2)+1)}
\frac{x^{3/2}}{3/2}+\frac{x^{1/2}}{1/2}
\frac{2}{3} x^{3/2} + 2x^{1/2}